# FormulaQA: A Question Answering Dataset for Formula-Based Numerical Reasoning

### v2

@@ 3.3 Formula Database Construction
Our next step was to construct a unified formula repository for the entire dataset. Given that parameters in the same formula can be expressed differently across various problem contexts, for instance, the two formulas “[weight of water] = [mass of water] * [gravitational acceleration]” and “[weight] = [mass] * [gravitational acceleration]” both calculate the weight of an object, we need to merge these formulas into a single representation.
We divided the construction process of the formulas database into three steps: 1) Merge the formulas through symbolic rules. 2) Merge the formulas through semantics based method. 3) Manual review and error correction. In Table 3, we present the initial number of formulas and the remaining number of formulas after each step.
Step # Formulas
Before merging 12,906
After symbolic rules based merging 1,163
After semantics based merging 439
After manual review and error correction 272
@@
In this step, we merged formulas through symbolic rules. Specifically, this was achieved by comparing the structure of the formulas and the symbols contained within them. Taking this as an example of judging whether two formulas have the same structure: the formulas “$f_{1}:~{}a_{1}$=($b_{1}$+$c_{1}$)/$d_{1}$”, “$f_{2}:a_{2}$=($b_{2}$+$c_{2}$)/$d_{2}$” and “$f_{3}:b_{1}$=$a_{1}$*$d_{1}$-$c_{1}$” have the same structure because $f_{2}$ can be derived from $f_{1}$ by renaming parameters, and $f_{3}$ can be obtained from $f_{1}$ by moving $b_{1}$ to the left side of the equation. Moreover, in physics, certain physical quantities are conventionally represented by specific symbols. For example, the mass of an object is often denoted by “$m$” and the density of an object is frequently represented by the symbol “$\rho$”. Subscripts are then used to distinguish which specific object the physical quantity refers to, such as “$\rho_{water}$” for the density of water. For any two formulas, we first computed all transformations of a formula to obtain a set of all variants of the formula. Then, we compared the formula structures in the two sets to determine if two formulas were structurally equivalent. If they shared the same structure, we then compared whether the symbols, with subscripts removed, were identical. If they were, we considered these two formulas to be mergeable. When merging, we retained the parameter with the shorter length from the two.
Semantics based merging.
In the symbolic rules based merging process, the semantic information of the parameter names was neglected. This led us to perform merges grounded on the semantics of the parameter names. For instance, two formulas that were not merged during the symbol fusion stage, “[density] = [mass] / [volume]” and “[density of water ] = [mass of water] / [volume of water]”, can actually be merged. We would carry out the merging of these two formulas based on the semantic information of the parameter names (for example, "density" and "density of water" are semantically similar). For formulas with identical structures, we tokenized each pair of corresponding parameters to create two sets of words [^3] We used jieba: https://github.com/fxsjy/jieba.. When the two sets overlapped, the parameters were considered to have semantic connection, and the formulas became candidates for merging. Utilizing this approach, we identified a list of potential formula mergers and then consulted the LLM for a thorough evaluation of each pair. The prompts can be found in Appendix A.1.3.
@@
Upon completing the aforementioned merging process, we manually inspected the correctness of the merge, rectified instances where errors occurred in the merging, and manually merged formulas that were overlooked by the LLM.
@@ 4 Experiments Setup
In this section, we explore several methods to address the questions within FormulaQA. These include prompting LLMs using zero-shot and few-shot chains of thoughts (CoTs, Wei et al., 2022; Kojima et al., 2022), fine-tuning smaller models with less than 2 billion parameters, and training a formula retriever to retrieve formulas which are then incorporated into the LLM prompts.
@@ 4.1 Dataset Split
We divided FormulaQA into training, id (in-distribution) test, and ood (out-of-distribution) test, comprising 4,608, 421 and 391 questions, respectively. We required that all formulas in the id test must appear in the training set, whereas in the ood test, each question involves at least one formula that has not been seen in the training set. This division is designed to evaluate the generalization capabilities of fine-tuned models on formulas they have not previously encountered.
4.2 LLMs
@@
Following Kojima et al., 2022, we incorporated the phrase “Let’s think step by step” into the zero-shot prompt to guide LLMs in generating the reasoning steps. For the few-shot setting, we randomly sampled five questions from the train set to serve as examples for in-context learning. Each example includes the question text and the reasoning steps (i.e., the explanation). Examples of the prompts can be found in Appendix A.2.1.
@@
We conducted experiments on GPT-3.5-turbo, GLM4, and Qwen-max, with each of these models having over 100 billion parameters. We also evaluated on Llama2-7B Touvron et al. (2023), Qwen-7B/14B Bai et al. (2023), InternLM2-7B/20B Team (2023), including the base and chat versions of these models. We followed the common practice that few-shot experiments were performed on the base versions, while zero-shot experiments were conducted on the chat versions.
4.3 Fine-tuned Small Models
To examine the performance of smaller models with fine-tuning, we fine-tuned small models with sizes less than 2B parameters. We conducted experiments on mT5-base, mT5-large Xue et al. (2021), and Qwen-1.8B Bai et al. (2023) models. Given that small models have significantly weaker numerical calculation capabilities compared to LLMs Lin et al. (2020); Testolin (2023); Lu et al. (2023), we added a setting that allows the small models to perform numerical reasoning without conducting calculations themselves, instead relying on an external calculator. Specifically, one setting involves the small models directly computing the results, while the other setting requires the models to generate the formulas for the reasoning steps along with the corresponding numerical values and units of the parameters, which are then calculated with the aid of an external calculator.
4.4 Formula Retriever
@@
We trained a formula retriever targeting the formula database on the training set. Specifically, we encoded each question using the Chinese-BERT-wwm-base Devlin et al. (2019); Cui et al. (2021) model to obtain the CLS vector of the question. Each formula in the formula database was represented by a randomly initialized vector. During training, we calculated the cosine score between the question vector and the formula vector. The retriever was then trained with in-batch negatives and contrastive learning loss Gao et al. (2021). Subsequently, for each question in the id test, we retrieved the top five formulas with the highest scores and included them in the prompt to observe the change in the performance of the LLM when provided with relevant formulas. More details can be found in Appendix A.2.2.
4.5 Implementation Details
We accessed to GPT-3.5-turbo [^4] https://platform.openai.com/docs, GLM4555https://open.bigmodel.cn/, and Qwen-max [^6] https://help.aliyun.com/zh/dashscope/developer-reference/quick-start through API calls with the default hyper-parameters. For other LLMs, we conducted experiments on NVIDIA V100-32G GPUs for 7B models, and on NVIDIA A100-80G GPUs for 14B/20B models. These LLMs generated using nucleus sampling with top_p=0.8. Models that require fine-tuning were experimented on NVIDIA V100 GPUs with Huggingface Transformers and Pytorch 2.0. For mT5-base and mT5-large, we set a learning rate of 5e-5 and a batch size of 32, testing the model after training for 50 epochs. For Qwen-1.8B, we used a learning rate of 1e-5 and a batch size of 32, and tested the model after training for 10 epochs. For the smaller models, we also utilized nucleus sampling with top_p=0.8 for generation. In the case of the smaller models, which directly outputted formulas along with corresponding parameter values and units, if the generation output contained formatting errors, we allowed the small model to retry up to 5 times until a correctly formatted output was generated. Training mT5-base, mT5-large, and Qwen-1.8B models requires 6, 12, and 12 hours respectively.
@@ 4.6 Metric
We utilized the numbat tool to evaluate the predictions generated by the model against the gold-standard answers. A prediction is deemed correct if the relative error (prediction - gold) / gold is less than 1%. We employed accuracy, which is the proportion of questions answered correctly, as our metric.
@@ 5.1 Results of LLMs
Model Size id test ood test Avg. zero-shot CoTs GPT-3.5-turbo unknown 26.13 25.58 25.87 GLM4 >100B 65.32 65.22 65.27 Qwen-max >100B 58.67 57.80 58.25 InternLM-chat 20B 5.70 4.60 5.17 Qwen-chat 14B 32.07 37.60 34.73 Qwen-chat 7B 7.36 8.70 8.01 InternLM-chat 7B 7.84 7.67 7.76 Llama2-chat 7B 0.00 0.26 0.13 few-shot CoTs GPT-3.5-turbo unknown 32.07 29.92 31.03 GLM4 >100B 62.47 65.98 64.16 Qwen-max >100B 58.91 63.94 61.33 InternLM-base 20B 18.29 11.25 14.90 Qwen-base 14B 44.89 36.83 41.01 Qwen-base 7B 21.14 18.16 19.71 InternLM-base 7B 9.50 8.18 8.86 Llama2-base 7B 1.43 0.26 0.87
Table 4: Results of LLMs on FormulaQA.
@@
The evaluation results on LLMs are shown in Table 4. GLM4 exhibited the best performance in both zero-shot and few-shot settings, surpassing the second-ranked Qwen-max by an average of 7.02 points in both id test and ood test. GPT-3.5-turbo, in comparison to GLM4 and Qwen-max, demonstrated inferior performance, which could be attributed to the likelihood that GLM4 and Qwen-max had access to a larger corpus of Chinese data during training, thereby enhancing their capability in Chinese. Among models with size not exceeding 20B, Qwen-14B demonstrated commendable performance in both zero-shot and few-shot settings. The subpar performance of Llama2 might be due to its training data being primarily in English. We also conducted few-shot testing on the chat version of LLMs with size not exceeding 20B, and the results can be found in Appendix A.2.3.
After incorporating few-shot examples, both GPT-3.5-turbo and Qwen-max demonstrated performance improvements, ranging from 0.24 to 6.14. However, similar performance changes were not observed in GLM4, possibly due to supervised fine-tuning and alignment with human preferences, which enhanced GLM4’s understanding of instructions but may have compromised its in-context learning ability.
The evaluation results indicate that even LLMs larger than 100B parameters can only achieve a maximum score of 65.98 on FormulaQA and the performance on LLMs with size not exceeding 20B is particularly unsatisfactory, underscoring the challenging nature of FormulaQA as an unresolved dataset.
5.2 Results of Fine-tuned Small Models
Model Size id test ood test Avg. w/o calculator mT5-base 960M 1.43 5.12 3.21 mT5-large 1.7B 19.00 10.74 15.02 Qwen-1.8B 1.8B 18.29 12.28 15.40 w/ calculator mT5-base 960M 33.49 30.18 31.90 mT5-large 1.7B 33.74 32.27 33.03 Qwen-1.8B 1.8B 64.78 54.45 59.81
Table 5: Results of fine-tuned small models on FormulaQA.
Table 5 shows the evaluation results for the fine-tuned small models, with and without the calculator. Recalling Section 4.1, because small models struggle with direct numerical calculations, we provided a calculator for additional support during evaluation. Except for mT5-base, which scored poorly, mT5-large and Qwen-1.8B performed better on the id test compared to the ood test, both with and without a calculator, indicating that out-of-domain formulas indeed affected model performance. However, despite the decrease in performance on the ood test, we observed that the small model still maintains generalization ability in formula prediction. We hope that dividing our test set into id test and ood test will be helpful for evaluating the generalization ability of fine-tuned models in future works.
@@ 5.2 Results of Fine-tuned Small Models
After using calculators, the performance of the three small models improved by scores ranging from 18.01 to 44.41. Our findings revealed that focusing on reasoning while delegating numerical computations to the calculator could enhance the performance of small models, given their limited calculating capability. This discovery provides valuable insights for future deployment of numerical reasoning systems on small models.

### v3

@@ 3.3 Formula Database Construction
Step # Formulas Before merging 12,906 After symbolic rules based merging 1,163 After semantic-based merging 439 After manual review and error correction 272
@@
Our next step was to construct a unified formula database for the entire dataset. Given that parameters in the same formula can be expressed differently across various problem contexts, for instance, the two formulas “[weight of water] = [mass of water] * [gravitational acceleration]” and “[weight] = [mass] * [gravitational acceleration]” both calculate the weight of an object, we need to merge these formulas into a single representation.
We divided the construction process of the formula database into three steps: 1) Merge the formulas through symbolic rules. 2) Merge the formulas through semantic-based method. 3) Manual review and error correction. In Table 3, we present the initial number of formulas and the remaining number of formulas after each step.
@@
In this step, we merged formulas through symbolic rules. Specifically, this was achieved by comparing the structure of the formulas and the symbols.Take the following as an example of judging whether two formulas have the same structure: the formulas “$f_{1}:~{}a_{1}$=($b_{1}$+$c_{1}$)/$d_{1}$”, “$f_{2}:a_{2}$=($b_{2}$+$c_{2}$)/$d_{2}$” and “$f_{3}:b_{1}$=$a_{1}$*$d_{1}$-$c_{1}$” have the same structure because $f_{2}$ can be derived from $f_{1}$ by renaming parameters, and $f_{3}$ can be obtained from $f_{1}$ by transformation. Moreover, in physics, certain physical quantities are conventionally represented by specific symbols. For example, the mass of an object is often denoted by “$m$” and the density of an object is frequently represented by the symbol “$\rho$”. Subscripts are then used to distinguish which specific object a physical quantity refers to, such as “$\rho_{water}$” for the density of water. For any two formulas, we first computed all the transformations of each formula to obtain a set of all its variants. Then, we compared the formula structures in the two sets to determine if two formulas were structurally equivalent. If they shared the same structure, we then compared whether their symbols, with subscripts removed, were identical. If they were, we considered these two formulas to be mergeable. When merging, we retained the parameter with the shorter length from the two. After merging based on symbolic rules, we reduced the number of formulas in the formula database from 12,906 to 1,163.
Semantic-based merging.
In the symbolic rules based merging process, the semantic information of the parameter names was neglected. This led us to perform merges grounded on the semantics of the parameter names. For instance, two formulas that were not merged during the symbolic fusion stage, “[density] = [mass] / [volume]” and “[density of water ] = [mass of water] / [volume of water]”, can actually be merged. We would carry out the merging of these two formulas based on the semantic information of the parameter names (for example, "density" and "density of water" are semantically similar). Specifically, for formulas with identical structures, we tokenized each pair of corresponding parameters to create two sets of words [^4] We used jieba: https://github.com/fxsjy/jieba.. When the two sets overlapped, the parameters were considered to have semantic connection, and the formulas became candidates for merging. Utilizing this approach, we identified a set of pairs of potentially mergeable formulas and then consulted the LLM for a thorough evaluation of each pair. The prompts can be found in Appendix A.1.3. After this step, the number of formulas in the formula database was reduced to 439.
@@
Upon completing the aforementioned merging process, we manually inspected the correctness of the results, rectified instances where errors occurred during merging, and manually merged formulas that were overlooked by the LLM. In this process, there were two human volunteers cross-validating the results of manual review and annotation. Finally, we obtained a formula database consisting of 272 formulas.
@@ 4 Experiments Setup
In this section, we explore several methods for handling the questions within FormulaReasoning, including prompting LLMs using zero-shot and few-shot chain-of-thought (CoT, Wei et al., 2022; Kojima et al., 2022), and training a formula retriever to retrieve formulas to be incorporated into LLM prompts. Additionally, we employed two approaches to enhancing the reasoning abilities of fine-tuned models with fewer than 7B parameters. The first approach involved dividing the reasoning process into distinct steps: formula generation, parameter extraction, and numerical calculation. The second approach leveraged data augmentation to improve the models’ reasoning ability.
@@ 4.1 Dataset Split
We divided FormulaReasoning into into subsets for training, id (in-distribution) test, and ood (out-of-distribution) test, comprising 4,608, 421 and 391 questions, respectively. We required that all formulas in the id test must appear in the training set, whereas in the ood test, each question involves at least one formula that has not been seen in the training set. This division is designed to evaluate the generalizability of fine-tuned models on formulas that they have not previously encountered.
4.2 Evaluation
4.2.1 Human Performance
We recruited 108 students from a high school, with each student being assigned 7–8 questions. Each student was given 40 minutes to complete these questions. These questions were used as part of their in-class exercises, and at the end, each student received a gift. The final statistics were collected to evaluate human performance, which was consented by all the students.
4.2.2 LLMs
@@
Following Kojima et al., 2022, we incorporated the phrase “Let’s think step by step” into the zero-shot prompt to guide LLMs in generating the reasoning steps. For the few-shot setting, we randomly sampled five questions from the training set to serve as examples for in-context learning. Each example includes the question text and the reasoning steps (i.e., the explanation). Examples of the prompts can be found in Appendix A.4.1.
@@
We conducted experiments on GPT-4-turbo, GPT-3.5-turbo, GLM4, and Qwen-max, with each of these models having over 100 billion parameters. We also evaluated on Llama2-7B (Touvron et al., 2023), Llama3-8B (Meta, 2024), Qwen-7B/14B (Bai et al., 2023), InternLM2-7B/20B (Team, 2023), ChatGLM3-6B (Zeng et al., 2022), including the base and chat versions of these models. We followed the common practice that few-shot experiments were performed on the base versions, while zero-shot experiments were conducted on the chat or instruct versions.
4.2.3 Formula Retriever
@@
We trained a formula retriever on the training set. Specifically, we encoded each question using the Chinese-BERT-wwm-base (Devlin et al., 2019; Cui et al., 2021) model to obtain the CLS vector of the question. Each formula in the formula database was represented by a randomly initialized vector. During training, we calculated the cosine score between the question vector and the formula vector. The retriever was then trained with in-batch negatives and contrastive learning loss (Gao et al., 2021). Subsequently, for each question in the id test, we retrieved the top five formulas with the highest scores and included them in the prompt to observe the change in the performance of the LLM when provided with relevant formulas. More details can be found in Appendix A.4.2.
4.2.4 Supervised Fine-tuned Models
We found that directly prompting models possessing fewer than 7B parameters failed to produce satisfactory outcomes (for example, ChatGLM3-6B attained merely 8.99 points in a zero-shot setting). Therefore, we conducted supervised fine-tuning of models with fewer than 7B parameters, yet discerned that, dissimilar to larger models (such as GPT-4-turbo), smaller models did not exhibit proficient performance in numerical extraction and calculation. In order to augment the reasoning capabilities of smaller models, we explored two approaches for improvement.
Chain-of-Thought Supervised Fine-Tuning (CoT-SFT)
We decomposed the reasoning process into several steps. First, we instructed the model to generate the formulas required to solve the question. Subsequently, the parameter names within the formulas were extracted, allowing the model to retrieve the corresponding values and units from the context. Next, the formulas and the associated parameter values were provided to a calculator to obtain the final result. This approach relieved the model of the numerical calculation, allowing it to concentrate on the reasoning aspect.
Data Augmentation (DA)
We augmented the training dataset with the assistance of larger models. Firstly, we utilized a few-shot approach to prompt the LLM (Qwen-max) to generate new question-answer pairs. The correctness of the computation process generated by the LLM was meticulously verified using a calculator. Subsequently, the formulas generated by the model were extracted and normalized. More details could be found in Appendix A.3.1.
@@ 4.3 Metric
We utilized numbat to evaluate the predictions generated by the model against the gold-standard answers. A prediction is deemed correct if the relative error (prediction - gold) / gold is less than 1%. We employed accuracy, which is the proportion of questions answered correctly, as our metric.
@@ 5 Experiments Results
In this section, we presented the experimental results and analysis. Due to space constraints, the error analysis can be found in Appendix A.2 and the implementation details can be found in Appendix A.4.
5.1 Human Performance
In FormulaReasoning, humans achieved impressive performance, with a score of 93.49 on the id test, 90.47 on the ood test, and an average score of 92.03.
@@ 5.2 Results of LLMs
Model Size zero-shot CoT few-shot CoT id test ood test Avg. id test ood test Avg. GPT-4-turbo unknown 70.07 72.89 71.43 71.50 77.49 74.38 GPT-3.5-turbo unknown 26.13 25.58 25.87 32.07 29.92 31.03 GLM4 >100B 65.32 65.22 65.27 62.47 65.98 64.16 Qwen-max >100B 58.67 57.80 58.25 58.91 63.94 61.33 InternLM 20B 5.70 4.60 5.17 18.29 11.25 14.90 Qwen 14B 32.07 37.60 34.73 44.89 36.83 41.01 Llama3 8B 26.66 17.98 20.41 12.81 8.87 10.91 Llama2 7B 0.00 0.26 0.13 1.43 0.26 0.87 Qwen 7B 7.36 8.70 8.01 21.14 18.16 19.71 InternLM 7B 7.84 7.67 7.76 9.50 8.18 8.86 ChatGLM3 6B 9.36 8.62 8.99 23.89 19.95 21.92 Human - 93.49 90.47 92.03 93.49 90.47 92.03
Table 4: Results of LLMs with zero-shot and few-shot prompting. indicates that the chat or instruct version of the model was used in the zero-shot setting, while the base version of the model was used in the few-shot setting.
@@
The evaluation results on LLMs are shown in Table 4. GPT-4-turbo exhibited the best performance in both zero-shot and few-shot settings, surpassing the second-ranked GLM4 by an average of 6.16 points in zero-shot setting and 10.22 in few-shot setting. Among models with size not exceeding 20B, Qwen-14B demonstrated commendable performance in both zero-shot and few-shot settings. The subpar performance of Llama2 might be due to its pre-training data being primarily in English. We also conducted few-shot testing on the chat version of LLMs with size not exceeding 20B, and the results can be found in Appendix A.4.3.  After incorporating few-shot examples, GPT-4-turbo, GPT-3.5-turbo and Qwen-max demonstrated performance improvements, ranging from 0.24 to 6.14. However, similar performance changes were not observed on GLM4, possibly due to its supervised fine-tuning and alignment with human preferences which enhanced GLM4’s understanding of instructions but probably also compromised its in-context learning ability.
Human performance surpassed the performance of few-shot GTP-4-turbo on the id and ood tests by margins of 21.99 and 13.25 points, respectively. Such results demonstrated that there remained a substantial gap between the current capabilities of state-of-the-art LLMs and human performance. This was even more pronounced when considering smaller-scale models. These findings underscored the challenging nature of FormulaReasoning as an unresolved dataset, and that there was significant room for improvement in LLMs as they struggled to match human levels of reasoning.
5.3 Results of LLMs with Formula Retriever
Model zero-shot few-shot GLM4 65.32 62.47    + formula retriever 70.31 65.80 GPT-4-turbo 70.07 71.50    + formula retriever 68.17 67.00
Table 5: Results of LLMs with Formula Retriever on the id test.
The results of LLMs utilizing the formula retriever are shown in Table 5. We found that the impact on performance varied among different LLMs when incorporating retrieved formulas into prompts. We observed a positive enhancement on GLM4, with score increments of 4.99 and 3.33 with zero-shot and few-shot, respectively, on the id test. However, we observed a performance decline with GPT-4-turbo. Specifically, we found that the top 5 retrieved formulas often included irrelevant ones, as the number of formulas required varies for each problem. The presence of these extraneous formulas affected the model’s performance, indicating that there is considerable room for further research in utilizing a formula database.
5.4 Results of Supervised Fine-tuned Models
Model Size id test ood test Avg. Qwen-1.8B 1.8B 55.91 44.58 50.25    + DA 56.16 45.32 50.74    + CoT-SFT 73.65 74.38 74.00 ChatGLM-6B 6B 52.95 40.64 47.02    + DA 53.44 45.32 49.53    + CoT-SFT 74.63 73.89 74.23
Table 6: Results of supervised fine-tuned models on FormulaReasoning.
Table 6 shows the results for the supervised fine-tuned models, with and without CoT-SFT and DA, which were detailed in Section 4.2.4. In most settings, both models achieved higher scores on the id test than the ood test, yet they still exhibited considerable performance on the ood test. This indicates that 1) the ood formulas indeed impacted model performance and 2) the models still demonstrate generalizability. We hope that the division of id test and ood test will be helpful for assessing the generalization ability of fine-tuned models in future works.
@@ 5.4 Results of Supervised Fine-tuned Models
It was noteworthy that with CoT-SFT, Qwen-1.8B and ChatGLM3-6B, with a mere parameter count of 1.8B and 6B, respectively, achieved performance comparable to GPT-4-turbo (though such a comparison may not be entirely fair). This indicated that the incorporation of CoT-SFT and the use of calculators could significantly enhance the reasoning capabilities of small models. Our findings revealed that focusing on reasoning with CoT while delegating numerical calculation to a calculator could enhance the performance of small models, given their limited calculating capability. The assistance of LLMs for data augmentation could also enhance smaller models’ reasoning capability. This discovery provides valuable insights for future deployment of numerical reasoning systems on small models.
6 Conclusion and Limitations
We introduced FormulaReasoning, a dataset for formula-based numerical reasoning. We annotated the reasoning steps with formulas for each question with both manual and LLM-assisted efforts. Furthermore, we constructed a formula database after merging formulas with similar meanings, serving as an external knowledge base for subsequent retrieval-based/augmented approaches.  We evaluated FormulaReasoning across various sizes of LLMs, supervised fine-tuned models, and retrieval-augmented LLMs, demonstrating its challenging nature as an unresolved task. Our findings indicate substantial room for improvement of existing models on formula-based numerical reasoning, thus motivating future research efforts.
We have also translated the dataset into English unitizing LLMs. However, we have not yet accurately assessed the quality of the translated dataset. At present, we have not released the English version of the dataset, but we will do so later after ensuring the quality of the English dataset. Additionally, our dataset is limited to the domain of physics. Although junior high school physics is not overly complex and can be understood by most people, it is still possible to explore formula-based question answering data in other domains.
Acknowledgments and Disclosure of Funding
This work was supported by the CIPSC-SMP-Zhipu.AI Large Model Cross-Disciplinary Fund. We are grateful to Chao Li for his suggestions and efforts in the annotation.
Appendix A Appendix
A.1 Dataset Construction
A.1.1 Prompts in Formula Normalization
The process of formula normalization is delineated into three distinct stages: the generation of natural language explanations, the extraction of the associated parameters from the explanations, and the subsequent error correction phase. The initial two stages are illustrated in Figures 3 and 4. The third stage is further splited into three specific error categories, each addressed by a dedicated prompt: input errors, where the parameters mentioned in the explanation are absent from the question; calculation errors, which occur when the calculator reports an error during the computation process; and output errors, where the final computed answer is incorrect. We provide an example here focusing on prompts for correcting calculation errors, while prompts for the other two error types can be found in our code submission. The prompts designed to correct calculation errors are depicted in Figure 5. The entire normalization procedure employs a 6-shot prompting, an instance of which is provided herein for illustrative purposes.
A.1.2 Examples of Deleted Questions
A.1.3 Semantic-based Merging for Formula Database Construction
Semantic-based merging primarily employs the LLM to comprehend formulas, ascertain if two formulas are semantically equivalent, and subsequently determine whether they can be merged into a single formula. The prompt for this procedure is illustrated in Figure 7. This approach ensures that the nuanced meanings embedded within formulas are accurately captured and evaluated for potential merging, thereby enhancing the quality of formula database.
A.2 Case Study and Error Analysis